ICNNAI-2010 will be made up of the following three categories of session.

• Plenary Sessions

⋆ We already recieved the confirmation from the following two prestigious keynote guests

· Professor Shun-ich Amari

(Riken, Japan) · Professor Xin Yao

(University of Birmingham)

• Special Sessions

⋆ We now call for Proposals.

⋆ Please Visit at:

http://icnnai.bstu.by/call-for-special-sessions.html

• Regular Sessions

⋆ See the right-hand page of this brochure.

Besides these sessions, we plan other social events.

• Welcome Cocktai Party

⋆ 31 May 2010 from 18:00 (The evening one day before the opening.)

• Gala dinner

⋆ 2 June 2010 (After the session of the day.)

• Excursion

⋆ Bus tour to the forrest which they say biggest in Europe.

Belarus is a land of fairly well-kept secret. In this lovely wonder land, with luck, you’ll find something that could not be discovered elsewhere.

Regular Session will be as follows though the submission procedure will be open on 1 January 2010 (6 month before the conference.)

• Submission

⋆ We call for a substantially new research which has not been published previously

⋆ For submission guideline, please visit at

http://icnnai.bstu.by/paperformat.html

where you will find the template

(both in Word and LaTeX)

⋆ Submission should be made via the automatic submission and review system via the Easy-Chair See the conference web-page

http://icnnai.bstu.by/call-for-paper.html

• Blind Review

⋆ All papers received will be reviewed by at least two referees in the area of the topic. We will follow a blind reviewing policy. Therefore, the paper must not contain any references that would identify the authors.

• Date

⋆ Inportant dates we currently planned are:

· Submission due: 7 March 2010 · Notification of acceptance or rejection: 1 April 2010 · Camera ready due: 21 April 2010

• Registration fee

⋆ 125 Euro